C=500+0.4Yd, MPS = 0.6, I = 500, G = 700, X = 200 and M=50+0.2Y the Tax rate is 20% of Y. The population is 90 people. Calculate NI equilibrium


1
Expert's answer
2021-09-23T09:13:16-0400

S+T+M=I+G+X

"S = - Co + MPS\times Y"

"-500+0.6 Y+0.2Y+50+0.2Y=500+700+200"

Y=1850